What is the acceleration of applied net force of 10 N to a given mass of 20 kg?


1
Expert's answer
2021-02-12T06:11:15-0500
"F=ma,""a=\dfrac{F}{m}=\dfrac{10\ N}{20\ kg}=0.5\ \dfrac{m}{s^2}."